OpenStudy (anonymous):

Add (Reduce to lowest terms) 1/7 + 5/14

jhonyy9 (jhonyy9):

1/2

OpenStudy (rogue):

Multiply 1/7 by 2/2 so both fractions have the same denominator. Then simply add and reduce. 2/14 + 5/14 = 7/14 = 1/2